Xenon Company incurred $1,000,000 in research and development costs over an 18 month period to develop product X1. Sales of product X1 will begin at the end of the 18 month period. The sales price of product X1 will be $20 per unit. Fixed costs are $140,000 per month. Xenon’s expected return on sales (margin after fixed and variable costs) target is 25%. Sales volume is projected to be 20,000 units per month.

a. Compute the target cost per unit and the variable cost allowable for Xenon to achieve its profitability goal.

b. Compute Break Even Time for product X1.

c. Was this a good opportunity to pursue if you knew the product life cycle would be 24 months? Why or why not?
Answer #1

Research & Development Cost = $1000000

Fixed Cost per month = $ 140000

Sale Volume per month = 20000 Units

Sale Price @ $ 20 per unit

Profit Margin @ 25% of Sales

Target Cost per unit = 20-20@25%= $15 per unit

Fixed Cost per unit per month = 140000/20000 = $7 per unit

Variable Cost per unit = $15-$7=$8

Profit per unit per month = $20-$15=$5 after meeting Fixed and Variable Cost

Total Profit in a month = 20000*5=$100000

Break Even Time for Product X1 = Research & Development Cost/Total Profit in a month

= 1000000/100000

= 10 months

Yes, this a good opportunity to purse even if the product life cycle would be 24 months because the break even time for Product X1 is 10 months and thereafter there would be pure profit of $ 100000 per month for the balance 14 months.